OpenStudy (anonymous):

Find the derivative y' by implicit differentiation if xy=sin(pi*y). Find the slope of the tangent line to xy=sin(pi*y) at point (x,y)=(0,1)

OpenStudy (freckles):

did you find derivative of both sides yet?

OpenStudy (freckles):

xy is a product so use the product rule to find (xy)'

OpenStudy (anonymous):

so far i got that y'=cos(pi*y), but im not sure if thats right

OpenStudy (freckles):

I'm not sure how you got that answer. Anyways I'm asking you to find (xy)' use the product rule since you have a product

OpenStudy (anonymous):

would it be y+xy'

OpenStudy (freckles):

right (xy)'=y+xy'

OpenStudy (freckles):

now we need to find (sin(pi*y))'

OpenStudy (freckles):

you need the chain rule and constant multiply rule here

OpenStudy (freckles):

what is (pi *y)' ?

OpenStudy (anonymous):

just y+pi*y'?

OpenStudy (freckles):

looks like you tried to use the product rule but did so incorrectly

OpenStudy (anonymous):

wait no y

OpenStudy (anonymous):

opps um cos(pi*y)*y'

OpenStudy (freckles):

You don't need the product rule to find derivative of pi*y like you have a constant times a function derivative will be just that constant times the derivative of that function this is what the constant multiple rule says also we could do the product rule but watch what happens \[\frac{d(\pi y)}{dx}=\pi \frac{dy}{dx}+y \frac{d \pi}{dx}=\pi \frac{dy}{dx}+y(0)=\pi \frac{dy}{dx}=\pi y'\]

OpenStudy (freckles):

\[\frac{d(\sin(\pi y))}{dx}=\pi y' \cos(\pi y) \]

OpenStudy (freckles):

so you have both sides after the differentiating step

OpenStudy (freckles):

now you need to solve for y'

OpenStudy (anonymous):

quick question, pi isnt a constant?

OpenStudy (freckles):

pi is totally a constant

OpenStudy (freckles):

pi is always 3.14159....

OpenStudy (freckles):

it never changes value

OpenStudy (freckles):

that is why it is called a constant because it constantly remains the same value forever and ever

OpenStudy (anonymous):

so when you use the chain rule, since pi is a constant how come pi is still included in the answer. Because isn't the derivative of a constant 0?

OpenStudy (freckles):

did you see what happened when i used the product rule? also do you not know the constant multiple rule?

OpenStudy (freckles):

I copy and pasted what I said above \[\frac{d(\pi y)}{dx}=\pi \frac{dy}{dx}+y \frac{d \pi}{dx}=\pi \frac{dy}{dx}+y(0)=\pi \frac{dy}{dx}=\pi y'\] The only reason I used the product rule here is because you tried to use it on this and failed

OpenStudy (freckles):

you could have said well pi is a constant and since it is being multiplied to a function I could use the constant multiple rule which says (cy)'=cy' where c is a constant

OpenStudy (freckles):

\[(\pi y)'=\pi y'\]

OpenStudy (freckles):

also you the derivative of a constant is zero that is why i put dpi/dx=0 when using the product rule

OpenStudy (freckles):